# Does LLM Focus on the Right Words? Diagnosing Language Bias in LLM-based Recommenders

### v1

@@ 4. Experiments
  • RQ3: Does GDRT mitigate language bias?
@@ 4.1.1. Datasets
Three widely used real-world datasets—Amazon Toys and Games, Amazon Clothing, Shoes and Jewelry, and Amazon Books—are employed in our experiments. These datasets are commonly adopted in studies on LLM-based recommenders (Cui et al., 2024; Bao et al., 2025; Cao et al., 2024; Lee et al., 2024; Li et al., 2023a; Wang et al., 2025; Lee et al., 2024). To ensure fair comparison, we follow the data preprocessing procedures employed in recent literature (Bao et al., 2025; Cui et al., 2024; Wang et al., 2025; Bao et al., 2024). Specifically, we first apply the 5-core setting to the raw datasets. For user interaction sequences exceeding 11 interactions, we segment the sequences using a sliding window of length 11. The segmented sequences are then sorted in ascending order by timestamp and partitioned into training, validation, and test sets in an 8:1:1 ratio. Due to the large size of Amazon Books, we randomly retain 100,000 items prior to 5-core processing. The statistics of the processed datasets are summarized in Table 1.
@@ 4.1.2. Baselines
  • LLM-based RS (SFT (Bao et al., 2025), CFT (Zhang et al., 2024b), MSL (Wang et al., 2025), LLaRA (Liao et al., 2024b), A-LLM (Kim et al., 2024)): SFT introduces instruction-tuning templates to fine-tune LLMs on RS datasets. CFT incorporates a causal loss to strengthen the behavior sequence modeling capabilities of LLMs. MSL improves the loss function specifically tailoring it to optimize recommendation-oriented objectives. LLaRA incorporates embeddings from traditional models to enhance the ability of LLMs to capture collaborative signals. A-LLM further aligns these embeddings with their corresponding textual information.
  • Debiasing Methods for LLM-based RS (Reweight (Jiang et al., 2024), SPRec (Gao et al., 2025a), D3 (Bao et al., 2024)): Reweight addresses popularity bias by balancing recommendations using pre-calculated item weights. SPRec develops negative sampling techniques based on Direct Preference Optimization (DPO) (Rafailov et al., 2024) to reduce popularity bias. D3 mitigates length bias through enhanced decoding strategies.
@@ 4.1.3. Implementation Details
For all LLM-based methods, we adopt LLaMA3-3B (Dubey et al., 2024) as the backbone, with the number of training epochs set to 5. For inference, we employ Constrained Beam Search (CBS), following prior work (Wang et al., 2025; Bao et al., 2024), to ensure that the recommended items are drawn from the item set. The beam size is fixed at 10. For evaluation, we evaluate the NDCG@5 on the validation set for each epoch’s checkpoint and select the checkpoint with the highest score for testing. The corresponding results on the test set are reported as the final performance. For the hyperparameters in GDRT, the number of groups $G$ is tuned from $\{2,5,10\}$, and $\tau$ is tuned from $\{0.1,0.2,0.3,0.5,1.0\}$. To ensure fair comparisons, we utilize the source code provided by the original authors and tune the hyperparameters of all baseline methods according to the guidelines specified in their respective publications.
Figure 7. Ratio of attribution values between auxiliary tokens and user interaction tokens using SFT and GDRT. Left: task description vs. user interaction tokens. Right: prefix tokens of predicted item (take the first token) vs. user interaction tokens.:
@@ 4.2. Performance Comparison (RQ1 & RQ2)
Table 2 presents a comparative analysis of the proposed GDRT method against the baselines. Overall, GDRT demonstrates substantial performance improvements across all datasets. By mitigating language bias, reducing excessive reliance on auxiliary tokens, and enhancing the model’s ability to capture user-specific behavioral patterns, GDRT delivers marked gains in both accuracy and fairness. Specifically, GDRT achieves an average improvement of 24.29% in NDCG@10 compared with SOTA methods. In terms of fairness, GDRT significantly alleviates the model’s tendency to recommend items strongly correlated with auxiliary information, leading to an average reduction of 37.43% in MGU@5 relative to the SOTA baselines. In contrast, other comparison methods yield only marginal improvements, and their lack of explicit mechanisms to address language bias results in limited fairness enhancement.
@@
Figure 8. Distribution of Top‑1 recommended items from GDRT-trained LLMs across five groups defined by auxiliary-token relevance (Group 1: highest relevance, Group 5: lowest relevance), with test-set distribution for comparison.:
Figure 9. Proportion of Top‑1 recommendations belonging to Item Group 1 (most relevant with auxiliary tokens) during training with SFT and GDRT.:
@@ 4.3. In-depth Analysis (RQ3)
In this section, we present an empirical analysis of the effectiveness of GDRT in mitigating language bias. As shown in Figure 7, GDRT considerably reduces the ratio of attribution values assigned to auxiliary tokens relative to interaction tokens, compared with SFT, indicating that GDRT effectively alleviates the problem of over-reliance on auxiliary tokens. Furthermore, GDRT markedly mitigates the unfairness induced by language bias. As shown in Figure 8, the recommendation distributions generated by GDRT more closely align with the item distribution in the test set. In addition, Figure 9 reveals that GDRT alleviates the issue observed in SFT, where the training process continuously amplifies the recommendation of items highly correlated with auxiliary tokens.
@@ 4.4. Hyper-Parameter Sensitivities (RQ4)
Figure 10. Hyperparameter sensitivity analysis on group number $G$ (dashed: SFT baseline).:
Figure 11. Hyperparameter sensitivity analysis on parameter $\tau$ in DRO (dashed: SFT baseline).:
@@ 5.1. Sequential Recommendation
Sequential recommendation focuses on predicting the next item a user will be interested in based on their historical interactions. Compared to collaborative filtering, sequential recommendation incorporates temporal information and places greater emphasis on capturing the evolving patterns of user interests. With the advancement of deep learning, numerous architectures based on deep neural networks have been introduced into the sequential recommendation. For example, GRU4Rec (Hidasi et al., 2015) employs RNNs, while Caser (Tang and Wang, 2018) utilizes CNNs to effectively capture long-term dependencies and modeling user interest patterns from historical behavior. More advanced models such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) leverage self-attention mechanisms (Vaswani, 2017), enabling the identification of the most relevant parts within the sequence. Due to the dynamic nature of data distributions as time evolves, DROS (Yang et al., 2023) introduces DRO (Rahimian and Mehrotra, 2019; Wu et al., 2023a) to further enhance the model’s robustness against distributional shifts caused by temporal changes. The readers may refer to the survey (Fang et al., 2020; Wang et al., 2019) for more details.
Figure 12. Performance comparisons in terms of both recommendation accuracy and efficiency.:
@@ 5.2. Biases in LLM-based Recommendation
Large Language Models (LLMs), with their powerful capabilities in comprehension, reasoning, and extensive knowledge (Dubey et al., 2024; Achiam et al., 2023; Guo et al., 2025), have been widely applied to recommendation systems (RS) (Wu et al., 2024). One prominent paradigm is LLM-based RS (Li et al., 2023c). In this paradigm, LLMs are directly employed as the recommendation backbone, with prompts used to instruct the models to perform recommendation tasks. However, due to the discrepancy between the pre-training objectives of LLMs and the goals of recommendation systems, subsequent studies have explored fine-tuning LLMs on domain-specific recommendation datasets to enhance their task-specific capabilities (Li et al., 2023c; Bao et al., 2025, 2023; Zhu et al., 2024; Wang et al., 2025, 2024; Liao et al., 2024b; Kim et al., 2024; Chen et al., 2024).
Recent studies have extensively explored bias and fairness issues in LLM-based RS, such as popularity bias (Gao et al., 2025b; Lu et al., 2025; Jiang et al., 2024; Gao et al., 2025a; Liao et al., 2024a), position bias (Chen et al., 2023; Ma et al., 2023; Hou et al., 2024; Bito et al., 2025; Jiang et al., 2025; Dai et al., 2024; Luo et al., 2025; Zhang et al., 2024a), length bias (Bao et al., 2024), and bias stemming from LLMs’ preferences for specific item attributes (Li et al., 2023b; Deldjoo, 2024; Jiang et al., 2024; Zhang et al., 2025). Popularity bias refers to the model’s tendency to disproportionately recommend popular items. Mitigation strategies typically reweight samples at the item (Jiang et al., 2024) or token level (Gao et al., 2025b; Lu et al., 2025) to prevent over-optimization toward popular items, or increasing the negative sampling probability for such items based on DPO (Gao et al., 2025a; Liao et al., 2024a). Position bias reflects the model’s sensitivity to the ordering of candidate items and is commonly addressed by incorporating position-invariant priors during training (Luo et al., 2025; Zhang et al., 2024a) or inference (Hou et al., 2024; Bito et al., 2025; Ma et al., 2023) stage. Length bias refers to the tendency of LLMs to favor items with longer token sequences during inference. D3 (Bao et al., 2024) alleviates this issue by removing length normalization from the score computation process. Additionally, LLMs have been shown to exhibit unequal preferences for items with specific attributes (e.g., genre, creator identity, or recency) (Deldjoo, 2024). To address this issue, some methods employ prompt-engineering techniques to encourage fairer recommendation behavior (Li et al., 2023b), whereas others reweight items across attribute categories to alleviate attribute-level bias (Jiang et al., 2024; Zhang et al., 2025).
Nevertheless, existing research has largely overlooked language bias, which arises during fine-tuning and reflects the inherent over-reliance of LLMs on auxiliary tokens. Since existing debiasing methods fail to account for this factor, their effectiveness remains limited. CFT (Zhang et al., 2024b) seeks to enhance the modeling of users’ historical interactions. However, it suffers from notable limitations, including objective misalignment, difficulties in weight selection, and high computational overhead, which constrain its effectiveness. These limitations are examined in detail in Section 2.2.4.
@@ 5.3. Group Distributionally Robust Optimization
Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) is an optimization framework that operates over predefined sample groups, aiming to achieve consistent and reliable performance across them by emphasizing the optimizing of the worst-performing group during training. It has been widely applied in various domains (Sagawa et al., 2019; Zhao et al., 2023; Oren et al., 2019) and has demonstrated strong effectiveness in mitigating group disparities (Hashimoto et al., 2018; Zhao et al., 2023) and reducing models’ reliance on shortcut correlations (Han et al., 2024; Creager et al., 2021). Several studies have applied Group DRO to RS. For example, S-DRO (Wen et al., 2022) is employed to improve user experience for less-represented user groups that prefer interacting with unpopular items. PDRO (Zhao et al., 2023) boosts the representation learning of both sparse users and items without sacrificing the performance for popular ones.
@@ 6. Conclusion
In this work, we identify a key limitation of supervised fine-tuning (SFT) in LLM-based recommenders: it often induces Language Bias, whereby the model over-relies on auxiliary tokens (e.g., task descriptions and prefix-generated tokens) while underutilizing core user interaction information. This bias undermines recommendation accuracy and raises unfairness concerns. To address this issue, we introduce Group Distributionally Robust Optimization-based Tuning (GDRT), which aims to reduce the model’s over-reliance on auxiliary tokens by applying Group DRO across token groups with varying degrees of relevance to auxiliary tokens. Extensive experiments on multiple public datasets demonstrate that GDRT effectively mitigates language bias, thereby significantly improving recommendation accuracy and enhancing fairness.
@@ Appendix A Appendices
A.1. The proof of Lemma 3.1

### v2

@@ 4. Experiments
  • RQ3: Does GDRT mitigate context bias?
@@ 4.1.1. Datasets
Three widely used real-world datasets—Amazon Toys and Games, Amazon Clothing, Shoes and Jewelry, and Amazon Books [^3] https://jmcauley.ucsd.edu/data/amazon/index_2014.html—are employed in our experiments. These datasets are commonly adopted in studies on LLM-based recommenders (Cui et al., 2024; Bao et al., 2025; Cao et al., 2024; Lee et al., 2024; Li et al., 2023a; Wang et al., 2025b; Lee et al., 2024). To ensure fair comparison, we follow the data preprocessing procedures employed in recent literature (Bao et al., 2025; Cui et al., 2024; Wang et al., 2025b; Bao et al., 2024). Specifically, we first apply the 5-core setting to the raw datasets. For user interaction sequences exceeding 11 interactions, we segment the sequences using a sliding window of length 11. The segmented sequences are then sorted in ascending order by timestamp and partitioned into training, validation, and test sets in an 8:1:1 ratio. Due to the large size of Amazon Books, we randomly retain 100,000 items prior to 5-core processing. The statistics of the processed datasets are summarized in Table 1.
@@ 4.1.2. Baselines
  • LLM-based RS (SFT (Bao et al., 2025), CFT (Zhang et al., 2024b), MSL (Wang et al., 2025b), LLaRA (Liao et al., 2024b), A-LLM (Kim et al., 2024)): SFT introduces instruction-tuning templates to fine-tune LLMs on RS datasets. CFT incorporates a causal loss to strengthen the behavior sequence modeling capabilities of LLMs. MSL improves the loss function specifically tailoring it to optimize recommendation-oriented objectives. LLaRA incorporates embeddings from traditional models to enhance the ability of LLMs to capture collaborative signals. A-LLM further aligns these embeddings with their corresponding textual information.
  • Debiasing Methods for LLM-based RS (Reweight (Jiang et al., 2024), SPRec (Gao et al., 2025a), D3 (Bao et al., 2024)): Reweight addresses popularity bias by balancing recommendations using pre-calculated item weights. SPRec develops negative sampling techniques based on Direct Preference Optimization (DPO) (Rafailov et al., 2024) to reduce popularity bias. D3 mitigates amplification bias through enhanced decoding strategies.
@@ 4.1.3. Implementation Details
For all LLM-based methods, we adopt LLaMA3.2-3B (Dubey et al., 2024) as the backbone, with the number of training epochs set to 5. The prompt design follows (Bao et al., 2025; Wang et al., 2025b). For inference, we employ Constrained Beam Search (CBS), following prior work (Wang et al., 2025b; Bao et al., 2024), to ensure that the recommended items are drawn from the item set. The beam size is fixed at 10. For evaluation, we evaluate the NDCG@5 on the validation set for each epoch’s checkpoint and select the checkpoint with the highest score for testing. The corresponding results on the test set are reported as the final performance. For the hyperparameters in GDRT, the number of groups $G$ is tuned from $\{2,5,10\}$, and $\tau$ is tuned from $\{0.1,0.2,0.3,0.5,1.0\}$. To ensure fair comparisons, we utilize the source code provided by the original authors and tune the hyperparameters of all baseline methods according to the guidelines specified in their respective publications.
Figure 7. Ratio of attribution values between auxiliary tokens and user interaction tokens using SFT and GDRT. Left: task description vs. user interaction tokens. Right: prefix tokens of predicted item (take the first token) vs. user interaction tokens.
@@ 4.2. Performance Comparison (RQ1 & RQ2)
Table 2 presents a comparative analysis of the proposed GDRT method against the baselines. Overall, GDRT demonstrates substantial performance improvements across all datasets. By mitigating context bias, reducing excessive reliance on auxiliary tokens, and enhancing the model’s ability to capture user-specific behavioral patterns, GDRT delivers marked gains in both accuracy and fairness. Specifically, GDRT achieves an average improvement of 24.29% in NDCG@10 compared with SOTA methods. In terms of fairness, GDRT significantly alleviates the model’s tendency to recommend items strongly correlated with auxiliary information, leading to an average reduction of 37.43% in MGU@5 relative to the SOTA baselines. In contrast, other comparison methods yield only marginal improvements, and their lack of explicit mechanisms to address context bias results in limited fairness enhancement.
@@
Figure 8. Distribution of Top‑1 recommended items from GDRT-trained LLMs across five groups defined by auxiliary-token relevance (Group 1: highest relevance, Group 5: lowest relevance), with test-set distribution for comparison.
Figure 9. Proportion of Top‑1 recommendations belonging to Item Group 1 (highest relevance with auxiliary tokens) during training with SFT and GDRT.
@@ 4.3. In-depth Analysis (RQ3)
In this section, we present an empirical analysis of the effectiveness of GDRT in mitigating context bias. As shown in Figure 7, GDRT considerably reduces the ratio of attribution values assigned to auxiliary tokens relative to interaction tokens, compared with SFT, indicating that GDRT effectively alleviates the problem of over-reliance on auxiliary tokens. Furthermore, GDRT markedly mitigates the unfairness induced by context bias. As shown in Figure 8, the recommendation distributions generated by GDRT more closely align with the item distribution in the test set. In addition, Figure 9 reveals that GDRT alleviates the issue observed in SFT, where the training process continuously amplifies the recommendation of items highly correlated with auxiliary tokens.
@@ 4.4. Hyper-Parameter Sensitivities (RQ4)
Figure 10. Hyperparameter sensitivity analysis on group number $G$ (dashed: SFT baseline).
Figure 11. Hyperparameter sensitivity analysis on parameter $\tau$ in DRO (dashed: SFT baseline).
@@ 5.1. Sequential Recommendation
Sequential recommendation focuses on predicting the next item a user will be interested in based on their historical interactions. Compared to collaborative filtering (Chen et al., 2025a, 2024a; Yang et al., 2024; Zhang et al., 2025a; Yang et al., 2025; Wang et al., 2024a), sequential recommendation incorporates temporal information and places greater emphasis on capturing the evolving patterns of user interests. With the advancement of deep learning, numerous architectures based on deep neural networks have been introduced into the sequential recommendation. For example, GRU4Rec (Hidasi et al., 2015) employs RNNs, while Caser (Tang and Wang, 2018) utilizes CNNs to effectively capture long-term dependencies and modeling user interest patterns from historical behavior. More advanced models such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) leverage self-attention mechanisms (Vaswani, 2017), enabling the identification of the most relevant parts within the sequence. Due to the dynamic nature of data distributions as time evolves (Wang et al., 2024c, d), DROS (Yang et al., 2023) introduces DRO (Rahimian and Mehrotra, 2019; Wu et al., 2023a) to further enhance the model’s robustness against distributional shifts caused by temporal changes. The readers may refer to the survey (Fang et al., 2020; Wang et al., 2019) for more details.
Figure 12. Performance comparisons in terms of both recommendation accuracy and efficiency.
@@ 5.2. Biases in LLM-based Recommendation
Large Language Models (LLMs), with their powerful capabilities in comprehension, reasoning, and extensive knowledge (Dubey et al., 2024; Achiam et al., 2023; Guo et al., 2025; Team et al., 2024), have been widely applied to recommendation systems (Wu et al., 2024; Wang et al., 2025c; Cui et al., 2025b, a). One prominent paradigm is LLM-based RS (Li et al., 2023c), which directly leverages LLMs as the backbone of the recommender. Subsequent studies have explored fine-tuning LLMs on domain-specific recommendation datasets to enhance their recommendation capabilities (Li et al., 2023c; Bao et al., 2025, 2023; Zhu et al., 2024; Wang et al., 2025b, 2024b; Liao et al., 2024b; Kim et al., 2024; Chen et al., 2024b).
Recent studies have extensively explored bias and fairness issues in LLM-based RS, such as popularity bias (Gao et al., 2025b; Lu et al., 2025; Jiang et al., 2024; Gao et al., 2025a; Liao et al., 2024a; Lin et al., 2025), position bias (Chen et al., 2023; Ma et al., 2023; Hou et al., 2024; Bito et al., 2025; Jiang et al., 2025; Dai et al., 2024; Luo et al., 2025; Zhang et al., 2024a), amplification bias (Bao et al., 2024), and bias stemming from LLMs’ preferences for specific item attributes (Li et al., 2023b; Deldjoo, 2024; Jiang et al., 2024; Zhang et al., 2025b). Nevertheless, existing research has largely overlooked context bias, which arises during fine-tuning and reflects the inherent over-reliance of LLMs on auxiliary tokens. Since existing debiasing methods fail to account for this factor, their effectiveness remains limited. CFT (Zhang et al., 2024b) seeks to enhance the modeling of users’ historical interactions, but suffers from objective misalignment, weight selection challenges, and high computational cost, restricting its applicability. These limitations are further discussed in Section 2.2.4.
@@ 5.3. Group Distributionally Robust Optimization
Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) is an optimization framework that operates over predefined sample groups, aiming to achieve consistent and reliable performance across them by emphasizing the optimizing of the worst-performing group during training. It has been widely applied in various domains (Sagawa et al., 2019; Zhao et al., 2023; Oren et al., 2019) and has demonstrated strong effectiveness in mitigating group disparities (Hashimoto et al., 2018; Zhao et al., 2023) as well as reducing models’ reliance on shortcut correlations (Han et al., 2024; Creager et al., 2021). Several studies have also integrated Group DRO into RS to improve model performance for sparse users and items (Wen et al., 2022; Zhao et al., 2023).
@@ 6. Conclusion
In this work, we identify a key limitation of supervised fine-tuning (SFT) in LLM-based recommenders: it often induces Context Bias, whereby the model over-relies on auxiliary tokens (e.g., task descriptions and prefix-generated tokens) while underutilizing core user interaction information. This bias undermines recommendation accuracy and raises unfairness concerns. To address this issue, we introduce Group Distributionally Robust Optimization-based Tuning (GDRT), which aims to reduce the model’s over-reliance on auxiliary tokens by applying Group DRO across token groups with varying degrees of relevance to auxiliary tokens. Extensive experiments on multiple public datasets demonstrate that GDRT effectively mitigates context bias, thereby significantly improving recommendation accuracy and enhancing fairness.
@@ Appendix A Appendices
A.1. Verification of Context Bias Across Diverse Prompt Templates and LLMs
In this section, we extend the FAA experiments described in Section 2.2.1 to a broader range of prompt templates (see Figure 13) and LLM backbones (see Figure 14), in order to assess whether context bias is a widespread phenomenon. For prompt templates, we adopt those proposed in (Liao et al., 2024b) and (Bao et al., 2024), referred to as Prompt1 and Prompt2. For LLM backbones, we evaluate LLaMA3‑8B (Dubey et al., 2024) and Qwen2.5‑1.5B (Team et al., 2024). Across all prompt and backbone configurations, the results consistently reveal pronounced context bias, wherein SFT leads the model to over-rely on auxiliary tokens.
Figure 13. Ratio of attribution values between auxiliary tokens and user-interaction tokens before and after SFT across different prompt templates.
Figure 14. Ratio of attribution values between auxiliary tokens and user-interaction tokens before and after SFT across different LLMs.
A.2. The proof of Lemma 3.1